# IG Captioner: Information Gain Captioners are Strong Zero-shot Classifiers

### v1

@@ Abstract
**footnotetext: Work done while an intern at Google DeepMind.
@@ 2 Related Work
Classifier-Free Guidance (CFG). CFG [10] is an important approach to improve the sample quality for the text-to-image generative models [17, 23, 27, 7, 34]. During training, CFG requires the models to generate images both w/ and w/o the text conditions. During inference, the output image is sampled according to a linear combination of both the conditional and unconditional predictions to improve the text-image alignment, which relates CFG and IG captioner from a high level. Differently, IG captioner is an image-to-text captioner and uses the probability gain from unconditionally to conditionally generating the captions with the aim of reducing the bias from text priors for zero-shot classification tasks.
@@ 3 Method
3.1 Evaluating Captioners on Discriminative Tasks
Maximum Likelihood Estimation (MLE) Objective. Let $I\in\mathbb{R}^{H\times W\times 3}$, $T\in\mathbb{R}^{N\times D}$ and $t\in\mathbb{R}^{D}$ be the image, caption and the word in the caption, where $H$, $W$ are the height and width of the image and $N$, $D$ are the number of text tokens and the token dimension. A captioner is a function $F_{\theta}:I\rightarrow T$ where $\theta$ is the model parameter.
$\operatorname*{arg\,max}_{i}\left[\log{P(T_{i}|I)}\right]$ (1)
Low Performance Using MLE Objective. To study the effectiveness of MLE as objective, we pretrain a captioner on Laion-5B and use it to perform zero-shot ImageNet classification. We set up a strong baseline: a CLIP classifier which is pretrained with contrastive loss and tailored for classification tasks. The captioner has a image encoder and a text decoder while the CLIP classifier has a image encoder and a text encoder. Both of them have the same model size, training and evaluation settings to ensure fair comparisons. The results are shown in Tab. 1. When directly using the MLE objective (Eqn. (1)) the captioner shows $>19\%$ top-1 accuracy degradation compared with the CLIP classifier.
Result Analaysis of MLE Evaluation. Since the captioner is a vision-language model, we are interested in understanding how much visual information is provided by the visual inputs and utilized by the model. We pretrain a language model that has the same architecture as the text decoder of the captioner. It models the captions in the Laion-5B dataset and is able to predict $\log P(T)$. There are 1K classes on ImageNet and for each class we use the 80 prompts proposed by CLIP. In total, there are 80K unique captions. Given an image $I$, we consider $\log P(T|I)$ and $\log P(T)$ as two variables and compute their Pearson correlation coefficient (PCC) [19, 31] with 80K observations in  Tab. 2. We observe a strong correlation between $\log P(T|I)$ and $\log P(T)$, which indicates that the captioner is strongly biased by the text priors and tends to ignore the visual information. The correlation is also visualized in Fig. 3.
@@ 3.3 Training Information Gain Captioners
MSCOCO (5K test set) Flickr30K (1K test set)
image $\rightarrow$ text (%) text $\rightarrow$ image (%) image $\rightarrow$ text (%) text $\rightarrow$ image (%)
model encoder pretrain loss R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP classifier ViT-B contrast 56.3 80.3 87.5 38.5 64.7 74.6 85.2 97.2 98.7 68.2 88.9 93.6
captioner ViT-B generate 24.7 48.2 59.7 45.1 68.9 77.5 46.9 73.1 81.5 72.2 90.8 94.3
IG captioner ViT-B generate 47.4 73.7 82.5 42.9 67.4 76.1 60.6 90.0 94.7 70.0 89.3 93.6
CLIP classifier ViT-L contrast 59.3 83.0 89.5 42.1 67.9 77.4 88.0 98.3 99.0 71.1 90.8 94.7
captioner ViT-L generate 29.9 53.8 65.7 49.2 72.8 80.8 56.6 82.0 88.1 77.9 93.1 96.1
IG captioner ViT-L generate 52.8 77.9 86.0 48.5 71.6 80.2 64.7 93.4 97.7 76.2 92.6 95.6
Table 4: Zero-shot image-text retrieval on MSCOCO [2] and Flickr30K [20] dataset. All the models are pretrained on Laion-5B dataset.
4 Experiments
@@ 4.1 Zero-shot ImageNet classification
Settings. We use the English subset of the public Laion-5B [28] dataset as the pretraining dataset for all experiments. After pretraining, we perform zero-shot image classification on the ImageNet [5] validation dataset. There are 1000 classes in this validation set and each class has 50 images. We build the captioner and IG captioner with an image encoder and a text decoder, and build the CLIP [22] classifier with an image encoder and a text encoder. All the models consist of standard transformer blocks [32]. We set the layer number of image encoder and text decoder / text encoder the same. We use a batch size 4096 for all the experiments. During evaluation, we follow CLIP to use 80 prompts for each class. We use the voting mechanism to combine the predictions from different prompts for the captioner and IG captioner. All the training and evaluation procedures are the same to ensure fair comparisons.
Results. The results are summarized in Tab. 3. First, compared with standard captioners, IG captioner achieves 19.7% and 18.1% top-1 accuracy improvements when using ViT-B and ViT-L [6] as the image encoders. Second, compared with the strong discriminative baselines, CLIP classifiers, IG captioner achieves 0.5% / 1.6% top-1 accuracy improvements when using the image encoder ViT-B / ViT-L. These improvements demonstrate that IG captioner is a strong zero-shot classifier as a pure generative model. We measure the correlations of the evaluation objective and the caption prior in Sec. 5.1 to validate the effectiveness of the design of IG captioner. Besides, we perform the ablation studies in Sec. 5.3 to demonstrate the efficacy of the IG evaluation on the standard captioner.
@@ 4.2 Zero-shot Image-text Retrieval on MSCOCO and Flickr30K
Settings. We use the same models in Sec. 4.1 pretrined on Laion-5B to perform zero-shot image-text retrieval on MSCOCO 5K and Flickr30K 1K test sets. There are 5 captions for each image. In order to measure the recalls, we compute a similarity matrix $\mathbf{M}_{a\times b}$ where $a$ and $b$ are the number of images and captions. For the CLIP classifier, $\mathbf{M}_{i,j}=\text{E}_{I_{i}}\cdot\text{E}_{T_{j}}$, where $\text{E}_{T_{j}}$ and $\text{E}_{I_{i}}$ are the embeddings of the $i^{\text{th}}$ image and $j^{\text{th}}$ caption. For the captioner, we compute $\mathbf{M}_{i,j}=\log P(T_{j}|I_{i})$ while for the IG captioner, we compute $\mathbf{M}_{i,j}=\log P(T_{j}|I_{i})-\log P(T_{j})$.
Results on MSCOCO. We summarize the results on MSCOCO in Tab. 4. For the image $\rightarrow$ text retrieval, IG captioner significantly boosts the performance of a generative captioner. IG captioner shows 22.7% / 25.5% / 22.8% improvements over the captioner on R@1 / R@5 and R@10 with the ViT-B image encoder, and shows 22.9% / 24.1% / 20.3% imporvements with the ViT-L image encoder. Furthermore, IG captioner achieves comparable performances with the CLIP classifier. The performance gaps between the IG captioner and the CLIP classifier are 8.9% / 6.6% / 5.0% on R@1 / R@5 / R@10 with the ViT-B encoder. While with the ViT-L image encoder, the gaps are further decreased to 6.5% / 5.1% / 3.5% on R@1 / R@5 / R@10.
@@
Results on Flickr30K. The results are summarized in Tab. 4. We observe the similar phenomena on MSCOCO. For the image $\rightarrow$ text retrieval, on the one hand, IG captioner shows significant improvements over the captioner. With the ViT-B image encoder, IG captioner improves the R@1 / R@5 / R@10 by 13.7% / 16.9% / 13.2%, and with the ViT-L image encoder, the improvements are 8.1% / 11.4% / 9.6%. On the other hand, the performance gaps between the IG captioner and the CLIP classifier are significantly reduced. For R@1, the gaps are 24.6% / 23.3% with ViT-B / ViT-L image encoder. However, for R@5 / R@10, the gaps become 7.2% / 4.0% with ViT-B encoder, and 4.9% / 1.3% with ViT-L encoder. For the text $\rightarrow$ image retrieval, we observe the similar superiority of both the captioner and IG captioner over the CLIP classifier.
@@ 5.1 Weights of the Text Prior Removal and Correlation Measurements
approach encoder eval obj $\uparrow$ acc. (%)
@@
approach model model type pretrain data params top-1 acc. (%)
Text-to-Image Classifier [4] Imagen [26] text-to-image diffusion model L5B + Pty 2B 62.7
Diffusion Classifier [12] Stable Diffusion 2.0 [25] text-to-image diffusion model L5B 1.45B 61.4
IG captioner IG captioner w/ ViT-L image-to-text captioner L5B 1.07B 69.8
@@
Settings. In this section, we ablate the $\alpha$ in the evaluation objective of IG captioner, $\log P(T|I)-\alpha\log P(T)$. We use the same training and evaluation settings as the main experiments. We pretrain the IG captioner with the ViT-B encoder on the Laion-5B dataset and perform zero-shot ImageNet classification and zero-shot image-text retrieval on MSCOCO and Flickr30K. We ablation the $\alpha$ from $0.0$ to $1.0$ with a step size $0.1$. For the retrieval tasks, we do not list the text $\rightarrow$ image recalls because the value of $\alpha$ does not affect the predictions of text $\rightarrow$ image retrieval. In order to validate our goal of removing the bias from text priors, we report the Pearson Correlation Coefficient (PCC) of $\log P(T)$ and the evaluation objective when varying the value of $\alpha$.
Results. We summarize the results in Tab. 5. On the one hand, the smaller correlation between $\log P(T|I)$ and $\log P(T)$ generally corresponds to better performances. For the ImageNet classification, when $\alpha=0.08$, the best PCC 0.08 and performance 64.3% are achieved at the same time. For the retrieval tasks on MSCOCO / Flickr30K, the best PCC, 0.0 / 0.12, and the highest R@10, 82.5% / 95.8%, are obtained simultaneously with $\alpha=0.9$. On the other hand, it requires a degree of text priors to achieve the best fine-grained performances on specific datasets. The highest R@1, 47.4% / 63.4%, correspond to the PCC, 0.19 / 0.48, on MSCOCO / Flickr30K. Through this ablation studies, we set $\alpha=0.8$ for all the other experiments. $\alpha=0.0$ corresponds to the bad strong correlations, resulting in the poor performances. $\alpha=1.0$ causes the negative correlations and produces the unsatisfying performances.
5.2 Language Model + Captioner versus IG Captioner
Settings. IG captioner has a text decoder that is able to predict both $\log P(T|I)$ and $\log P(T)$. One variant of IG captioner consists of a pair of models. One of them is the unimodal Language Model (LM) to predict $\log P(T)$. The other one is the multimodal captioner, a standard captioner to predict $\log P(T|I)$. During inference, the prediction is based on the subtraction of their outputs. We call this LM + Cap approach.
Results. We report the results in Tab. 6. It is observed that the LM + Cap approach performs worse than the IG captioner. Besides, it is less efficient due to the extra language model. From another perspective, it demonstrates that the text decoder of IG captioner is able to perform both the multimodal and unimodal tasks well.
@@ 5.3 IG Evaluation
Settings. IG captioner differs from a standard captioner in both the training and evaluation phases. In this section, we study whether the IG evaluation will improve a standard captioner as a classifier. The IG evaluation requires the prediction of $\log P(T)$. Since the standard captioner can not directly predict this, we feed the captioner zero-intensity images to compute $\log P(T|\mathbf{0})$ as the approximation.
Results. We show the results in Tab. 5. The IG evaluation is able to significantly improve the performance of the captioners. Specifically, with the ViT-B / ViT-L image encoder, IG evaluation increases the zero-shot top-1 ImageNet classification accuracy by $9.0\%$ / $7.9\%$. Adding the training part of IG captioner, the performances will be further improved by $10.7\%$ / $10.2\%$.
@@ 5.4 Comparisons to Text-to-image Generative Classifiers
Settings. There are two kinds of vision-language generative models: text-to-image generative models and image-to-text captioners. There have been recent works [12, 4] that successfully convert the large text-to-image diffusion models to the zero-shot generative classifiers. We compare IG captioner with these models.
Results. The results are summarized in Tab. 7. With the same or smaller amount of pretraining data, IG captioner demonstrates the best performances. Compared with the Text-to-Image Classifiers [4] and Diffusion Classifier [12] using Imagen [26] and Stable Diffusion [25], respectively, IG captioner shows $7.1\%$ / $8.4\%$ zero-shot top-1 accuracy improvements on ImageNet. Besides, IG captioner is a more compact model and only uses $53.5\%$ / $73.8\%$ of their parameters.

### v2

@@ 2 Related Work
Classifier-Free Guidance (CFG). CFG [17] is an important approach to improving the sample quality of text-to-image generative models [30, 37, 42, 11, 50]. During training, CFG requires the model to generate images both with and without text conditions. During inference, the output image is sampled according to a linear combination of both the conditional and unconditional predictions to improve text-image alignment, which relates CFG to IG captioner from a high level. In contrast, IG captioner is an image-to-text captioner that uses the probability gain from unconditionally to conditionally generating captions with the aim of reducing the bias from text priors for zero-shot classification tasks.
@@
Image-to-text Multimodal Language Models. Multimodal language models have become the dominant approaches for image captioning, including OFA [48], Flamingo [2], BLIP [22, 23], CoCa [49], LLaVA [28, 26], MiniGPT-4 [51], InstructBLIP [8], Kosmos [18, 33], Qwen-VL [3], VFC [12], LLaVA-NeXT [27, 20], Gemini [46, 39], and GPT-4 [1]. Differently, IG captioner demonstrates the potential of pure generative training on discriminative tasks using public datasets.
@@ 3 Method
3.1 Evaluating Captioners on Discriminative Tasks
3.1.1 Maximum Likelihood Estimation (MLE) Objective.
Let $I\in\mathbb{R}^{H\times W\times 3}$, $T\in\mathbb{R}^{N\times D}$ and $t\in\mathbb{R}^{D}$ be the image, caption and the word in the caption, where $H$, $W$ are the height and width of the image and $N$, $D$ are the number of text tokens and the token dimension. A captioner is a function $F_{\theta}:I\rightarrow T$ where $\theta$ is the model parameter.
$\operatorname*{arg\,max}_{i}\left[\log{P(T_{i}|I)}\right]$ (1)
3.1.2 Low Performance Using MLE Objective.
To study the effectiveness of MLE as objective, we pretrain a captioner on Laion-5B and use it to perform zero-shot ImageNet classification. We set up a strong baseline: a CLIP classifier which is pretrained with contrastive loss and tailored for classification tasks. The captioner has a image encoder and a text decoder while the CLIP classifier has a image encoder and a text encoder. Both of them have the same model size, training and evaluation settings to ensure fair comparisons. The results are shown in Tab. 1. When directly using the MLE objective (Eqn. (1)) the captioner shows $>19\%$ top-1 accuracy degradation compared with the CLIP classifier.
@@ 3.1 Evaluating Captioners on Discriminative Tasks
3.1.3 Result Analaysis of MLE Evaluation.
Since the captioner is a vision-language model, we are interested in understanding how much visual information is provided by the visual inputs and utilized by the model. We pretrain a language model that has the same architecture as the text decoder of the captioner. It models the captions in the Laion-5B dataset and is able to predict $\log P(T)$. There are 1K classes on ImageNet and for each class we use the 80 prompts proposed by CLIP. In total, there are 80K unique captions. Given an image $I$, we consider $\log P(T|I)$ and $\log P(T)$ as two variables and compute their Pearson correlation coefficient (PCC) [32, 13] with 80K observations in  Tab. 2. We observe a strong correlation between $\log P(T|I)$ and $\log P(T)$, which indicates that the captioner is strongly biased by the text priors and tends to ignore the visual information. The correlation is also visualized in Fig. 3.
@@ 4.1 Zero-shot ImageNet classification
4.1.1 Settings.
We use the English subset of the public Laion-5B [43] dataset as the pretraining dataset for all experiments. After pretraining, we perform zero-shot image classification on the ImageNet [9] validation dataset. There are 1000 classes in this validation set and each class has 50 images. We build the captioner and IG captioner with an image encoder and a text decoder, and build the CLIP [36] classifier with an image encoder and a text encoder. All the models consist of standard transformer blocks [47]. We set the layer number of image encoder and text decoder / text encoder the same. We use a batch size 4096 for all the experiments. During evaluation, we follow CLIP to use 80 prompts for each class. We use the voting mechanism to combine the predictions from different prompts for the captioner and IG captioner. All the training and evaluation procedures are the same to ensure fair comparisons.
4.1.2 Results.
The results are summarized in Tab. 3. First, compared with standard captioners, IG captioner achieves 19.7% and 18.1% top-1 accuracy improvements when using ViT-B and ViT-L [10] as the image encoders. Second, compared with the strong discriminative baselines, CLIP classifiers, IG captioner achieves 0.5% / 1.6% top-1 accuracy improvements when using the image encoder ViT-B / ViT-L. These improvements demonstrate that IG captioner is a strong zero-shot classifier as a pure generative model. We measure the correlations of the evaluation objective and the caption prior in Sec. 5.1 to validate the effectiveness of the design of IG captioner. Besides, we perform the ablation studies in Sec. 5.3 to demonstrate the efficacy of the IG evaluation on the standard captioner.
@@ 4.2 Zero-shot Image-text Retrieval on MSCOCO and Flickr30K
4.2.1 Settings.
We use the same models in Sec. 4.1 pretrined on Laion-5B to perform zero-shot image-text retrieval on MSCOCO 5K and Flickr30K 1K test sets. There are 5 captions for each image. In order to measure the recalls, we compute a similarity matrix $\mathbf{M}_{a\times b}$ where $a$ and $b$ are the number of images and captions. For the CLIP classifier, $\mathbf{M}_{i,j}=\text{E}_{I_{i}}\cdot\text{E}_{T_{j}}$, where $\text{E}_{T_{j}}$ and $\text{E}_{I_{i}}$ are the embeddings of the $i^{\text{th}}$ image and $j^{\text{th}}$ caption. For the captioner, we compute $\mathbf{M}_{i,j}=\log P(T_{j}|I_{i})$ while for the IG captioner, we compute $\mathbf{M}_{i,j}=\log P(T_{j}|I_{i})-\log P(T_{j})$.
4.2.2 Results on MSCOCO.
We summarize the results on MSCOCO in Tab. 4. For the image $\rightarrow$ text retrieval, IG captioner significantly boosts the performance of a generative captioner. IG captioner shows 22.7% / 25.5% / 22.8% improvements over the captioner on R@1 / R@5 and R@10 with the ViT-B image encoder, and shows 22.9% / 24.1% / 20.3% imporvements with the ViT-L image encoder. Furthermore, IG captioner achieves comparable performances with the CLIP classifier. The performance gaps between the IG captioner and the CLIP classifier are 8.9% / 6.6% / 5.0% on R@1 / R@5 / R@10 with the ViT-B encoder. While with the ViT-L image encoder, the gaps are further decreased to 6.5% / 5.1% / 3.5% on R@1 / R@5 / R@10.
@@
MSCOCO (5K test set)
image $\rightarrow$ text (%) text $\rightarrow$ image (%)
model encoder pretrain loss R@1 R@5 R@10 R@1 R@5 R@10
CLIP classifier ViT-B contrast 56.3 80.3 87.5 38.5 64.7 74.6
captioner ViT-B generate 24.7 48.2 59.7 45.1 68.9 77.5
IG captioner ViT-B generate 47.4 73.7 82.5 42.9 67.4 76.1
CLIP classifier ViT-L contrast 59.3 83.0 89.5 42.1 67.9 77.4
captioner ViT-L generate 29.9 53.8 65.7 49.2 72.8 80.8
IG captioner ViT-L generate 52.8 77.9 86.0 48.5 71.6 80.2
Table 4: Zero-shot image-text retrieval on MSCOCO [5]dataset. All the models are pretrained on Laion-5B dataset.
Flickr30K (1K test set)
image $\rightarrow$ text (%) text $\rightarrow$ image (%)
model encoder pretrain loss R@1 R@5 R@10 R@1 R@5 R@10
CLIP classifier ViT-B contrast 85.2 97.2 98.7 68.2 88.9 93.6
captioner ViT-B generate 46.9 73.1 81.5 72.2 90.8 94.3
IG captioner ViT-B generate 60.6 90.0 94.7 70.0 89.3 93.6
CLIP classifier ViT-L contrast 88.0 98.3 99.0 71.1 90.8 94.7
captioner ViT-L generate 56.6 82.0 88.1 77.9 93.1 96.1
IG captioner ViT-L generate 64.7 93.4 97.7 76.2 92.6 95.6
Table 5: Zero-shot image-text retrieval on Flickr30K [34] dataset. All the models are pretrained on Laion-5B dataset.
4.2.3 Results on Flickr30K.
The results are summarized in Tab. 4. We observe the similar phenomena on MSCOCO. For the image $\rightarrow$ text retrieval, on the one hand, IG captioner shows significant improvements over the captioner. With the ViT-B image encoder, IG captioner improves the R@1 / R@5 / R@10 by 13.7% / 16.9% / 13.2%, and with the ViT-L image encoder, the improvements are 8.1% / 11.4% / 9.6%. On the other hand, the performance gaps between the IG captioner and the CLIP classifier are significantly reduced. For R@1, the gaps are 24.6% / 23.3% with ViT-B / ViT-L image encoder. However, for R@5 / R@10, the gaps become 7.2% / 4.0% with ViT-B encoder, and 4.9% / 1.3% with ViT-L encoder. For the text $\rightarrow$ image retrieval, we observe the similar superiority of both the captioner and IG captioner over the CLIP classifier.
@@ 5.1 Weights of the Text Prior Removal and Correlation Measurements
5.1.1 Settings.
In this section, we ablate the $\alpha$ in the evaluation objective of IG captioner, $\log P(T|I)-\alpha\log P(T)$. We use the same training and evaluation settings as the main experiments. We pretrain the IG captioner with the ViT-B encoder on the Laion-5B dataset and perform zero-shot ImageNet classification and zero-shot image-text retrieval on MSCOCO and Flickr30K. We ablation the $\alpha$ from $0.0$ to $1.0$ with a step size $0.1$. For the retrieval tasks, we do not list the text $\rightarrow$ image recalls because the value of $\alpha$ does not affect the predictions of text $\rightarrow$ image retrieval. In order to validate our goal of removing the bias from text priors, we report the Pearson Correlation Coefficient (PCC) of $\log P(T)$ and the evaluation objective when varying the value of $\alpha$.
5.1.2 Results.
We summarize the results in Tab. 6. On the one hand, the smaller correlation between $\log P(T|I)$ and $\log P(T)$ generally corresponds to better performances. For the ImageNet classification, when $\alpha=0.08$, the best PCC 0.08 and performance 64.3% are achieved at the same time. For the retrieval tasks on MSCOCO / Flickr30K, the best PCC, 0.0 / 0.12, and the highest R@10, 82.5% / 95.8%, are obtained simultaneously with $\alpha=0.9$. On the other hand, it requires a degree of text priors to achieve the best fine-grained performances on specific datasets. The highest R@1, 47.4% / 63.4%, correspond to the PCC, 0.19 / 0.48, on MSCOCO / Flickr30K. $\alpha=0.0$ corresponds to the bad strong correlations, resulting in the poor performances. $\alpha=1.0$ causes the negative correlations and produces the unsatisfying performances. According to the ImageNet classification results with ViT-B encoder, we set $\alpha=0.8$ for all experiments in this paper. For the ImageNet classification with ViT-L encoder and all the image-text retrieval tasks, we do not change this $\alpha$.
5.2 Language Model + Captioner versus IG Captioner
5.2.1 Settings.
IG captioner has a text decoder that is able to predict both $\log P(T|I)$ and $\log P(T)$. One variant of IG captioner consists of a pair of models. One of them is the unimodal Language Model (LM) to predict $\log P(T)$. The other one is the multimodal captioner, a standard captioner to predict $\log P(T|I)$. During inference, the prediction is based on the subtraction of their outputs. We call this LM + Cap approach.
5.2.2 Results.
We report the results in Tab. 7. It is observed that the LM + Cap approach performs worse than the IG captioner. Besides, it is less efficient due to the extra language model. From another perspective, it demonstrates that the text decoder of IG captioner is able to perform both the multimodal and unimodal tasks well.
approach encoder eval obj $\uparrow$ top-1 acc. (%)
@@ 5.2 Language Model + Captioner versus IG Captioner
generative classifier model type pretrain data params top-1 acc. (%)
Imagen [7, 41] text-to-image L5B + Pty 2B 62.7
Stable Diffusion 2.0 [19, 40] text-to-image L5B 1.45B 61.4
IG captioner image-to-text L5B 1.07B 69.8
@@ 5.3 IG Evaluation
5.3.1 Settings.
IG captioner differs from a standard captioner in both the training and evaluation phases. In this section, we study whether the IG evaluation will improve a standard captioner as a classifier. The IG evaluation requires the prediction of $\log P(T)$. Since the standard captioner can not directly predict this, we feed the captioner zero-intensity images to compute $\log P(T|\mathbf{0})$ as the approximation.
5.3.2 Results.
We show the results in Tab. 5. The IG evaluation is able to significantly improve the performance of the captioners. Specifically, with the ViT-B / ViT-L image encoder, IG evaluation increases the zero-shot top-1 ImageNet classification accuracy by $9.0\%$ / $7.9\%$. Adding the training part of IG captioner, the performances will be further improved by $10.7\%$ / $10.2\%$.
@@ 5.4 Comparisons to Text-to-image Generative Classifiers
5.4.1 Settings.
There are two kinds of vision-language generative models: text-to-image generative models and image-to-text captioners. There have been recent works [19, 7] that successfully convert the large text-to-image diffusion models to the zero-shot generative classifiers. We compare IG captioner with these models.
5.4.2 Results.
The results are summarized in Tab. 8. With the same or smaller amount of pretraining data, IG captioner demonstrates the best performances. Compared with the Text-to-Image Classifiers [7] and Diffusion Classifier [19] using Imagen [41] and Stable Diffusion [40], respectively, IG captioner shows $7.1\%$ / $8.4\%$ zero-shot top-1 accuracy improvements on ImageNet. Besides, IG captioner is a more compact model and only uses $53.5\%$ / $73.8\%$ of their parameters.